I need help with the following problems. Solve each of the following for slope intercept form, then graph the line by plotting the y intercept and counting the slope.
1. 4x-2y=8
2. 3x+4y=12

Answer by stanbon(75887)

I need help with the following problems. Solve each of the following for slope intercept form, then graph the line by plotting the y intercept and counting the slope.
1. 4x-2y=8
y = 2x-4
slope = 2/1 = 2
y-intercept: (0,-4)
Plot (0,-4)
Move your pencil 1 to the right, then 2 up to (1,-2)
Draw a line thru those two points
======================================================
2. 3x+4y=12
y = (-3/4)x+4
slope = -3/4
intercept: (0,4)
Plot (0,4)
Move your pencil 4 to the right, then down 3 to (4,1)
Draw a line thru those two points.
=======================================
Cheers,
Stan H.